OpenStudy (anonymous):

the line through (1,0) and parallel to x-y=7 what is the slope intercept form?

OpenStudy (anonymous):

slope is same as slope of \[x-y=7\] \[y=x-7\] slope is 1 point slope formula give \[y-0=1(x-1)\] \[y=x-1\]

OpenStudy (anonymous):

slope of x-y=7 is 1, not -1

OpenStudy (anonymous):

yup.. u r right